Please I need help this one ASAP

3. "You want to design your office. For this purpose, your purchased office furniture for $37,000, paid $4,900 for its freight and $1200 for installation. What is the tax depreciation amount you can take in year 5? (Assume a 7-Year MACRS)"
Answer #1
Depreciation amount in year 5 $   3,848.83
Working:
Calculation of cost of Furniture:
Purchase price $       37,000
Freight $         4,900
Installation $         1,200
Cost of Furniture $       43,100
Depreciation Schedule:
Year Cost Depreciation rate Depreciation expense Accumulated Depreciation expense Book Value at the end of year
a b c=a*b d e=a-d
1 $       43,100 14.29% $   6,158.99 $   6,158.99 $ 36,941.01
2 $       43,100 24.49% $ 10,555.19 $ 16,714.18 $ 26,385.82
3 $       43,100 17.49% $   7,538.19 $ 24,252.37 $ 18,847.63
4 $       43,100 12.49% $   5,383.19 $ 29,635.56 $ 13,464.44
5 $       43,100 8.93% $   3,848.83 $ 33,484.39 $   9,615.61
6 $       43,100 8.92% $   3,844.52 $ 37,328.91 $   5,771.09
7 $       43,100 8.93% $   3,848.83 $ 41,177.74 $   1,922.26
8 $       43,100 4.46% $   1,922.26 $ 43,100.00 $ 0.00
Total 100.00% $ 43,100.00